Ship reports banner with links Home IODP-USIO daily/weekly reports

< November | Next expedition >

IODP Expedition 329

South Pacific Gyre Microbiology

Daily Science Report for 12-13 December 2010

LOCATION: Wynyard Wharf, Auckland, New Zealand

SCIENCE UPDATE: Our transit home, laboratory cleaning and shipment preparations continued through 12 December. We arrived at the pilot station at 0500 hr on 13 December and were berthed at the Wynyard Wharf in Auckland, New Zealand before 0700 hr, ending Expedition 329.


Daily Science Report for 11 December 2010

LOCATION: Underway to Auckland, New Zealand. Midnight position: 38° 6.7'S 179° 47.3'W

SCIENCE UPDATE: Our transit home continues. All Expedition Reports have been finalized and turned in to the Publications Specialist/Yeoperson. Laboratory cleaning and shipment preparations are ongoing. The end of cruise meeting was held to discuss the post-expedition meetings. Estimated time of arrival at Auckland pilot station is 0500 hr on 13 December.


Daily Science Report for 10 December 2010

LOCATION: Underway to Auckland, New Zealand. Midnight position: 39° 56.6'S 175° 43.1'W

SCIENCE UPDATE: Our transit home continues. All cores from Site U1371 have been processed for petrophysical properties and described. Reports from Site U1371 have been submitted to the co-chiefs' office for review and the Expedition Summary report is being assembled. Scientists met to present the scientific results from Site U1371. Data box shipments are being prepared. Laboratory cleaning and disassembling and packing of third party instrumentation are ongoing. Estimated time of arrival at Auckland pilot station is 0600 hr on 13 December. Clocks were advanced 24 hrs when crossing the dateline.


Daily Science Report for 8 December 2010

LOCATION: Underway to Auckland, New Zealand. Midnight position: 42° 7.3'S 171° 32.2'W

SCIENCE UPDATE: Our transit home continues under rough sea and heavy swell. We continue to process and analyze core samples from Site U1371 in the Core, Geochemistry and Microbiology, and Radio-isotope Laboratories. Writing and review of Sites U1370 and U1371 reports continues, as well as assembling parts of Expedition Summary report. Non-frozen/non-refrigerated sample shipments are being prepared and laboratory cleaning started. The expedition logo printing party took place in the conference room. Estimated time of arrival at Auckland pilot station is 0600 hr on 13 December.


Daily Science Report for 7 December 2010

LOCATION: Underway to Auckland, New Zealand. Midnight position: 43° 57.0'S 167° 31.7'W

SCIENCE UPDATE: Expedition scientists and USIO technical support staff continue to process core and samples from Site U1371 in the Core, Geochemistry and Microbiology Laboratories. Core sections from Hole U1371F are being measured for natural gamma, density, magnetic susceptibility, thermal conductivity, and P-wave velocity on the laboratory track systems. Up to Core U1371F-3H have been split into working and archived halves, digitally imaged and visually described. Complete archive sections are being measured in the cryogenic magnetometer. The broad range of biogeochemical analyses of samples from Site U1371 continued. Stable and radio isotope-tracer experiments for quantifying microbial activities and processes are ongoing in the radioisotope laboratory.


Daily Science Report for 6 December 2010

LOCATION: Underway to Auckland, New Zealand. Midnight position: 45° 58.4'S 163° 13.2'W

SCIENCE UPDATE: Coring operations at the last site of the expedition were completed. Fourteen cores were taken at Hole U1371F reaching a total penetration of 130.6 mbsf and recovering 118.44 m of sediment (90.6%). All cores from Hole U1371F were transferred directly from the catwalk to the core refrigerator and aseptically sampled for microbiological studies. Two additional holes were cored to provide more material of the upper part of the sedimentary column. We recovered a 1.37-m mudline core from Hole U1371G and a 6.2-m mudline core from Hole U1371H. These cores were transferred to the cold room in the Geochemistry and Microbiology Laboratory for oxygen measurements. The unsampled whole round section pieces from Holes U1371E through H are being measured for physical properties, split into working and archive halves, imaged and described. Biogeochemical and microbiological analyses of samples collected at Site U1371 continued. We started the final 1173 nautical mile transit to Auckland at 23:26 hrs.


Daily Science Report for 5 December 2010

LOCATION: Site U1371 (SPG-12A, 45° 57.8'S, 163° 11.1'W, water depth: 5304.8 m)

SCIENCE UPDATE: Cores U1371E-7H to 14H penetrated from 55.7 mbsf to basement depth at 128.2 mbsf. The total recovery for Hole U1371E was a 118.16 m (92 %). Hole U1371F, the microbiology assigned hole, was initiated and Cores U1371F-1H to 9H were recovered by midnight. Current penetration for this hole stands at 85.2 mbsf with a recovery of 72.21 m (84.7 %). Cores from Hole U1371F were transferred directly from the catwalk to the core refrigerator and aseptically sampled in high resolution for microbiological studies. All cores from Holes U1371D and U1371E were measured for density, magnetic susceptibility, natural gamma radiation, and P-wave velocity using the wholeround multisensor logger (WRMSL). The sediment in Hole U1371D ranges from clay bearing diatom ooze, to diatom bearing clay to zeolitic pelagic clay with increasing depth. Ash layers are present throughout the sequence. Samples from Hole U1371E are being measured for a broad range of biogeochemical species. Samples from Hole U1371F are being measured for microbial abundance and processed for shipboard and shorebased cultivation and molecular analyses. PFT testing of samples from Site U1371 is being conducted to check for potential contamination of drill water from the periphery toward the center of the cores and to confirm the suitability of core material for microbiological research.


Daily Science Report for 4 December 2010

LOCATION: Site U1371 (SPG-12A, 45° 57.8'S, 163° 11.1'W, water depth: 5304.8 m)

SCIENCE UPDATE: Cores U1371D-6H to 14H penetrated from 45.4 mbsf to basement (at 126 mbsf). A total of 126.9 meters of sediment (101%) were recovered from Hole U1371D. Samples for safety and analytical methane measurements were taken on the catwalk soon after core recovery. Cores were curated promptly and the sections then transferred to the cold room in the Geochemistry/Microbiology Laboratory for oxygen measurements. Coring operations were stopped for the second time at this site, after retrieving the last core from Hole U1371D, for the drill crew to change out a 20 ft knobby and the top saver sub due to a leak in the connection between them. Coring resumed after repairs were complete and the biogeochemistry-assigned Hole U1371E was initiated. Cores U1371E-1H to 6H (0 to 55.7 mbsf) were retrieved before midnight, recovering 47.77 m of sediment or 85.76 %. IW whole round core samples were taken on the catwalk for pore water chemistry. The curated core sections were then transferred to the hold deck’s core refrigerator for additional sampling for biogeochemical measurements. Cores U1371D-1H through 11H have been measured for physical properties in the Core Laboratory. The upper six cores from Hole U1371D have been split into working and archive halves, digitally imaged and described. The sediment consists of clay bearing diatom ooze with radiolarians. Review of Site U1370 reports continues.


Daily Science Report for 3 December 2010

LOCATION: Site U1371 (SPG-12A, 45° 57.8'S, 163° 11.1'W, water depth: 5304.8 m)

SCIENCE UPDATE: Four attempts were made to shoot the first mudline core at Site U1371 due to the combination of 4-5 meter heave conditions and water depth. The first core recovered from Hole U1371B came up empty, while the second one (U1371B-1H) was highly disturbed. A new hole was initiated and Core U1371C-1H recovered 9.83 m of sediment (103%). The completely filled core liner indicated that a good mudline core was not recovered. Consequently, the drill string was raised few meters and Hole U1371D was initiated obtaining a 7.4 m mudline core. Cores U1371D-1H to 5H penetrated to 45.4 mbsf and recovered 43.88 m of sediment (96.65% recovery). Subsequently, the forward coring line failed at the rope socket while pulling Core U1371D-6H. The APC barrel then fell ~3000 meters inside the hole. Coring operations were temporarily halted while the crew retrieved the coring line, the sinker bar assembly and the core barrel, and reheaded the coring line. APC coring in Hole U1371D resumed six hours later at 21:00 hrs. All cores recovered were taken to the cold room in the Microbiology Laboratory immediately after catwalk curation for oxygen measurements. Expedition scientists are also working on site reports.


Daily Science Report for 2 December 2010

LOCATION: Site U1371 (SPG-12A, 45° 57.8'S, 163° 11.1'W, water depth: 5304.8 m)

SCIENCE UPDATE: We arrived at Site U1371, the last site of Expedition 329, at 09:07 hrs. Site U1371 corresponds to Scientific Prospectus Site SPG-12A. The scientific objectives at this site are to determine the habitability, activity and microbial community composition in the subseafloor sediments just outside the gyre. In context with the other drill sites, the expedition scientists aim to resolve how different microbial communities are from the heart of the gyre to the southern edge and whether living microbes persist in this sediment for the last ~73 million years. In the ship laboratories, scientists continued to subsample core samples from Hole U1370F for onboard microbial cell enumeration, cultivation and shorebased molecular analyses. Isotope-tracer experiments for quantifying microbial activities and processes also continued. Hole U1371A was drilled without coring to confirm basement depth and the presence of any hard layers in the sedimentary sequence. Basement was confirmed at 5438.5 mbrf or 122.5 mbsl. A midnight meeting was held to present and discuss the scientific results from Site U1370 and the operations plan and scientific objectives for Site U1371.


Daily Science Report for 1 December 2010

LOCATION: Underway to Site U1371 (SPG-12A). Midnight position: 45° 15.4'S, 161° 23.9'W

SCIENCE UPDATE: The unsampled core sections from Hole U1370F (the microbiology-assigned hole) were brought to the Core Laboratory and integrated into the routine core flow. They were measured for physical properties using the whole-round and half-round multisensor track systems, imaged and visually described. Cores from Holes U1370E and U1370F consist of dark brown zeolotic metalliferous pelagic clay. The archive core section halves from Hole U1370D are being measured for paleomagnetic properties in the ship's cryogenic magnetometer. Samples from Hole U1370F are being prepared and processed for onboard cell enumeration, cultivation and shore-based molecular analyses. Stable and radio isotope-tracer experiments for quantifying microbial activities and processes are ongoing in the radioisotope van. Expedition scientists are also working on site reports.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Tuesday, 08-Nov-2011 08:46:00 CST.